# When Algorithms Meet Artists: Semantic Compression of Artists’ Concerns in the Public AI-Art Debate

### v3

@@ 4.1 Overview: What the Findings Imply
Our results provide strong evidence that public discourse on AI and art acts as a low-pass filter on stakeholder complexity. While artists articulate a high-dimensional landscape of 34 distinct frames spanning nuanced positions on threat, utility, ownership, transparency, and compensation Lovato and others [2024], public discourse compresses this variance into a narrow region of the semantic map. Across distributional, geometric, and stratified analyses, we find that artist stakeholder frames are not proportionally represented in the public discourse topic landscape, even after accounting for stylistic form.
This pattern is consistent with recent media research showing that public debates about AI grant disproportionate standing to economic and scientific actors while marginalizing ethical and social frames Zai and others [2025], and with arguments that agenda-setting is increasingly shaped by algorithmic curation and engagement-driven ranking rather than epistemic completeness Sichach [2025], Pane [2025]. Huang and Gadavanij Huang and Gadavanij [2025] document analogous dynamics in AI-in-education discourse, where affected stakeholders receive systematically less coverage than institutional actors despite the direct relevance of their concerns. The core empirical signature we document—95% of artist probes concentrated in 4 of 22 public topics—provides a measurable instantiation of what Baeyaert Baeyaert [2025] terms marginalization without explicit exclusion: affected-group knowledge is systematically backgrounded relative to institutional and technical narratives even when present in the discourse environment.
The magnitude of this compression is striking. Artists in our sample hold internally diverse and often conflicting positions: some embrace AI as a creative tool, while others view it as an existential threat; some would donate their work freely, while others demand ongoing royalties; some support mandatory transparency, while others remain uncertain about ownership claims Lovato and others [2024], Kawakami and Venkatagiri [2024]. Yet this heterogeneity—the “territory” of stakeholder reality—maps onto a small fraction of the public discourse “map.” The remaining 18 topics, comprising over 60% of public discourse volume, contain virtually no artist perspective.
@@ 4.2.1 Public Discourse Is Broad, But Salience Is Uneven
The public discourse landscape on AI art appears pluralistic at first. Our semantic map contains 22 topics spanning technical genealogies, institutional and market narratives, governance and law, philosophy of creativity, and tool practice. This breadth is consistent with recent topic-mapping work showing that public debates about image-generative AI contain multiple distinct interpretive clusters rather than a single linear narrative Saeedi and Taleghani [2025], Banks and Li [2025].
However, the key question is not whether multiple topics exist, but which topics become salient and which frames gain standing—a distinction central to framing theory Entman [1993]. Prior work suggests that AI discourse frequently foregrounds progress/economic potential and technologically oriented narratives while ethical-social concerns receive less attention or are simplified Zai and others [2025], Bøgh [2025]. Bøgh Bøgh [2025] specifically documents this pattern in AI-art media coverage, finding that nuanced labor concerns are reduced to “AI as Threat” versus “AI as Tool” framings. Our results extend this logic by showing that stakeholder-derived frames occupy a comparatively narrow region of the public topic space: 4 topics account for 95% of artists’ concerns, while 14 topics contain none whatsoever.
@@ 4.2.2 Agenda Misalignment Manifests as Semantic Compression
Importantly, this is not a claim that artists lack nuance. Lovato et al. Lovato and others [2024], Kawakami and Venkatagiri Kawakami and Venkatagiri [2024], and Jiang et al. Jiang and others [2023] document fine-grained and internally divergent positions. Instead, our contribution is to show that the public discourse “map” does not allocate proportional representational space to this stakeholder “territory.” In the public sphere, the distinction between an artist who would donate their work freely (the “Commons Donor” frame) and an artist who demands profit-sharing (the “Value Capture” frame) is erased; both land in Cluster 9. The 18 distinct ownership frames—ranging from exclusive artist rights to shared user-artist claims to rejection of all property frameworks—all collapse into Cluster 17. The public agenda has “captured” the issue of AI art but has stripped it of its internal political logic, reducing complex labor disputes to broad tropes of “resistance” or “democratization.”
@@ 4.3.1 What Occupies the Artist-Absent Regions?
If artists concentrate in 4 clusters, what populates the other 18? Our analysis of artist-absent clusters reveals a discourse landscape dominated by voices and concerns orthogonal to artist experience. This pattern is consistent with epistemic marginalization Baeyaert [2025], wherein affected-group knowledge is backgrounded through differential amplification rather than explicit exclusion.
@@
1. Institutional and market narratives (35.6% of public corpus): These topics feature curators, executives, and institutional voices discussing market implications, funding models, NFT economies, and the philosophical possibilities of AI art, without engaging the labor realities that concern practitioners. This pattern mirrors the actor asymmetries Huang and Gadavanij Huang and Gadavanij [2025] document in AI-education discourse, where institutional voices systematically overshadow those of affected practitioners.
2. Technical genealogies and spectacle (18.5% of public corpus): These topics focus on novelty, historical lineage, and technological milestones from DeepDream Mordvintsev et al. [2015] through GANs Goodfellow et al. [2014] to the Christie’s Belamy auction. This “origin story” discourse captures public attention through spectacle but does not address consent, compensation, or creative integrity.
@@
This pattern is consistent with the idea that the public sphere can operate as an exclusionary space, not through active silencing, but through differential amplification. By flooding discourse with technical, institutional, and spectacular narratives, the environment structurally backgrounds practitioners’ experiential knowledge Huang and Gadavanij [2025], Zai and others [2025], Baeyaert [2025].
@@ 4.3.2 Algorithmic Agenda-Setting and the “Post-Mediatized” Sphere
A distinctive feature of the contemporary discourse environment is that agenda-setting is increasingly algorithmic. Sichach Sichach [2025] argues that attention is now shaped not only by editorial institutions but also by recommender systems and algorithmic ranking that prioritize engagement and scale. Pane Pane [2025] similarly conceptualizes a “post-mediatized” sphere in which algorithmic curation helps set the epistemic boundaries of what is visible and discussable.
@@ 4.4.1 Threat and Utility Travel Well; Governance-Rights Concerns Do Not
One mechanism for differential salience is narrative portability. Media analyses of AI debates show recurring binary frames (e.g., opportunity versus crisis, tool versus threat) that simplify complexity into repeatable tropes Bøgh [2025], Cheung [2024]. These become dominant frames—interpretive packages that achieve salience through repetition and institutional amplification, crowding out alternative viewpoints.
@@
By contrast, governance-rights themes (transparency, ownership, compensation) map predominantly to Cluster 17, which receives only 5.84% of public discourse mass despite containing 40.6% of artist probes, a salience ratio of $6.95\times$. Governance-rights themes require technical pipeline details (e.g., how training data is sourced), legal doctrines (e.g., copyright, licensing, consent), and institutional enforcement mechanisms (e.g., opt-out registries, compensation schemes). These elements are more complex to compress into fast-moving media frames and thus more vulnerable to the compression we document Sichach [2025], Zai and others [2025], Pane [2025].
@@ 4.4.2 Coalition Mobilization Accentuates the Stakes
Recent organizing by creator coalitions underscores that transparency, consent, and compensation are not fringe concerns but actively contested governance claims. Over 100,000 visual artists in the UK have mobilized through the AOI Coalition against unauthorized use of copyrighted works in AI training AOI Coalition of Artists [2025]; the US Creators Coalition on AI, launched in December 2025, pursues similar objectives Creators Coalition on AI [2025]. The Andersen v. Stability AI class action Andersen et al. v. Stability AI et al. [2023] demonstrates that artist grievances have reached federal litigation, with plaintiffs alleging that generative models trained without consent constitute mass infringement.
@@ 4.7 Limitations
Methodological constraints: Artist probes are structured responses constrained by survey framing Lovato and others [2024], not organic discourse. The public corpus is shaped by Google Search retrieval, potentially overrepresenting highly visible sources. Our use of a single embedding model (e5-large-v2) may introduce representational biases.
@@ 4.8 Conclusion
Together, these findings constitute an empirical signature of epistemic marginalization Baeyaert [2025]—a structural backgrounding of affected-group knowledge that occurs without explicit silencing or malicious intent, through the combined operation of agenda-setting, algorithmic amplification, and milestone-driven spectacle Sichach [2025], Pane [2025], Cheung [2024].
@@ 5 Supplementary Results
5.1 Corpus and Map Details
5.1.1 Search Phrases
Table 2 presents the search phrases used to construct the public discourse corpus.
Table 2: Breadth of search phrases and prevalence of relevant sources. Left: search phrases used to build the public discourse corpus. Right: the number of relevant documents found in the first page of Google search results for each phrase. For some search terms, Google supplemented the standard 10 text links with up to 5 suggested video results, so the maximum possible total per search could reach up to 15 (values above 10 reflect this).
@@ 5.1.1 Search Phrases
5.1.2 Full Topic Inventory
Table 3 presents the complete 22-topic inventory with labels, keywords, and confidence scores from our multi-annotator interpretation process.
Table 3: Complete list of cluster labels, keywords, and final LLM judge’s confidence in interpretations.
@@ 14 GAN-era AI art explainer/history GANs, Obvious/Belamy (Christie’s), Robbie Barrat, latent space 8
5.1.3 Cluster Distribution by Data Source
Table 4 shows the distribution of public discourse chunks, public probes, and artist probes across all 22 topics. Bold values indicate notable concentrations.
Table 4: Cluster distribution by data source.
@@ 5.1.3 Cluster Distribution by Data Source
Figure 3: Two-dimensional UMAP projection of the public discourse semantic space showing all 22 topic clusters. Topics containing artist perspectives are circled in red (Topics 1, 8, 9, 10, 16, 17, 18). Artist concerns concentrate in a narrow region of the semantic space, while the majority of topics (particularly those in the upper portion of the map) contain no artist perspective.
5.2 Base Map Stability Validation
  • Naive coordinate averaging (31 seeds): Mean ARI =0.56=0.56
  • Distance-matrix consensus (31 seeds): Mean ARI =0.71=0.71
  • Relative improvement: $\sim$27%
5.3 Robustness Tests
5.3.1 Media Type and Year Associations
  • Media type $\times$ cluster: Cramér’s $V=0.453$
  • Year $\times$ cluster: Cramér’s $V=0.539$
  • Year $\times$ cluster within Articles: Cramér’s $V=0.542$
  • Year $\times$ cluster within Audio: Cramér’s $V=0.433$
  • Year $\times$ cluster within Papers: Cramér’s $V=0.710$
5.4 Semantic Pathway Analysis
Analysis of per-artist cross-theme patterns reveals:
  • Top 5 semantic pathways account for 73.4% of artists
  • Top 10 semantic pathways account for 89.4% of artists
Figure 4: Semantic pathway concentration analysis. The top 5 pathways account for 73.4% of all artists, and the top 10 pathways account for 89.4%, indicating stable cross-theme patterns rather than idiosyncratic scattering. Notably, governance concerns (Ownership and Transparency) exhibited the greatest internal diversity among the five concern dimensions—yet they are arguably the most compressed: Cluster 17 appears in all top 10 pathways for both Ownership and Transparency themes, collapsing this nuance into a single discursive region. By contrast, affective themes (Threat, Utility) distribute across Clusters 10 and 16, achieving slightly broader discursive footprint.
5.5 Additional Statistical Details
Full permutation test results for geometric separation:
  • Public vs. Artist Probes centroid distance: 2.034 ($p<0.0001$)
  • Public Probes vs. Artist Probes centroid distance: 1.336 ($p<0.0001$)
  • $k$-NN same-source rate (artist probes): 92.1% ($p<0.0001$)
  • $k$-NN same-source rate (public probes vs. artist): 76.5% ($p<0.0001$)
6 Supplementary Methods
6.1 Consensus UMAP Implementation
Figure 5: Consensus UMAP methodology. We generate 31 UMAP projections with different random seeds, compute pairwise distance matrices for each, average these matrices, and fit a final embedding from the consensus distance structure. This approach increases average seed-to-consensus Adjusted Rand Index from 0.56 (naive coordinate averaging) to 0.71 (distance-matrix consensus), yielding a more stable reference geometry.
6.1.1 Seed Pool and Multi-Run Generation
  • n_neighbors: 15
  • min_dist: 0.1
  • n_components: 8
  • metric: cosine
6.1.2 Distance-Matrix Consensus
6.1.3 UMAP Hyperparameter Selection
  • min_dist: 0.001 to 0.4 in evenly spaced steps
6.2 Projection Head Architecture
The MLP projection head consists of:
  • Input: 1024-dimensional e5-large-v2 embedding
  • Hidden layers: [512, 256, 128] with ReLU activation and dropout (0.2)
  • Output: 8-dimensional consensus UMAP coordinates
  • Training: Adam optimizer, MSE loss, 80/20 train/validation split
6.3 Use of Large Language Models
LLMs were used for the following tasks:
  1. 1.
    Anchor generation: GPT-5.1 generated 250 synthetic Likert-style statements
  2. 2.
    Validation filtering: Human + LLM pipeline retained on-theme public probes
  3. 3.
    Topic interpretation: GPT-5.1 Pro, Gemini Pro 1.5, and Claude Opus 4.5 independently labeled clusters; GPT-5.1 Pro served as consolidation judge
Human oversight was maintained throughout, and multiple models were used to mitigate single-model biases.

### v4

@@ 4.1 Overview: What the Findings Imply
Our results provide strong evidence that public discourse on AI and art acts as a low-pass filter on stakeholder complexity. While artists articulate a high-dimensional landscape of 34 distinct frames spanning nuanced positions on threat, utility, ownership, transparency, and compensation (Lovato et al., 2024), public discourse compresses this variance into a narrow region of the semantic map. Across distributional, geometric, and stratified analyses, we find that artist stakeholder frames are not proportionally represented in the public discourse topic landscape, even after accounting for stylistic form.
This pattern is consistent with recent media research showing that public debates about AI grant disproportionate standing to economic and scientific actors while marginalizing ethical and social frames (Zai et al., 2025), and with arguments that agenda-setting is increasingly shaped by algorithmic curation and engagement-driven ranking rather than epistemic completeness (Sichach, 2025; Pane, 2025). Huang and Gadavanij (2025) document analogous dynamics in AI-in-education discourse, where affected stakeholders receive systematically less coverage than institutional actors despite the direct relevance of their concerns. The core empirical signature we document—95% of artist probes concentrated in 4 of 22 public topics—provides a measurable instantiation of what Baeyaert (2025) terms marginalization without explicit exclusion: affected-group knowledge is systematically backgrounded relative to institutional and technical narratives even when present in the discourse environment.
The magnitude of this compression is striking. Artists in our sample hold internally diverse and often conflicting positions: some embrace AI as a creative tool, while others view it as an existential threat; some would donate their work freely, while others demand ongoing royalties; some support mandatory transparency, while others remain uncertain about ownership claims (Lovato et al., 2024; Kawakami and Venkatagiri, 2024). Yet this heterogeneity—the “territory” of stakeholder reality—maps onto a small fraction of the public discourse “map.” The remaining 18 topics, comprising over 60% of public discourse volume, contain virtually no artist perspective.
@@ 4.2.1 Public Discourse Is Broad, But Salience Is Uneven
The public discourse landscape on AI art appears pluralistic at first. Our semantic map contains 22 topics spanning technical genealogies, institutional and market narratives, governance and law, philosophy of creativity, and tool practice. This breadth is consistent with recent topic-mapping work showing that public debates about image-generative AI contain multiple distinct interpretive clusters rather than a single linear narrative (Saeedi and Taleghani, 2025; Banks and Li, 2025).
However, the key question is not whether multiple topics exist, but which topics become salient and which frames gain standing—a distinction central to framing theory (Entman, 1993). Prior work suggests that AI discourse frequently foregrounds progress/economic potential and technologically oriented narratives while ethical-social concerns receive less attention or are simplified (Zai et al., 2025; Bøgh, 2025). Bøgh (2025) specifically documents this pattern in AI-art media coverage, finding that nuanced labor concerns are reduced to “AI as Threat” versus “AI as Tool” framings. Our results extend this logic by showing that stakeholder-derived frames occupy a comparatively narrow region of the public topic space: 4 topics account for 95% of artists’ concerns, while 14 topics contain none whatsoever.
@@ 4.2.2 Agenda Misalignment Manifests as Semantic Compression
Importantly, this is not a claim that artists lack nuance. Lovato et al. (2024), Kawakami and Venkatagiri (2024), and Jiang et al. (2023) document fine-grained and internally divergent positions. Instead, our contribution is to show that the public discourse “map” does not allocate proportional representational space to this stakeholder “territory.” In the public sphere, the distinction between an artist who would donate their work freely (the “Commons Donor” frame) and an artist who demands profit-sharing (the “Value Capture” frame) is erased; both land in Cluster 9. The 18 distinct ownership frames—ranging from exclusive artist rights to shared user-artist claims to rejection of all property frameworks—all collapse into Cluster 17. The public agenda has “captured” the issue of AI art but has stripped it of its internal political logic, reducing complex labor disputes to broad tropes of “resistance” or “democratization.”
@@ 4.3.1 What Occupies the Artist-Absent Regions?
If artists concentrate in 4 clusters, what populates the other 18? Our analysis of artist-absent clusters reveals a discourse landscape dominated by voices and concerns orthogonal to artist experience. This pattern is consistent with epistemic marginalization (Baeyaert, 2025), wherein affected-group knowledge is backgrounded through differential amplification rather than explicit exclusion.
@@
1. Institutional and market narratives (35.6% of public corpus): These topics feature curators, executives, and institutional voices discussing market implications, funding models, NFT economies, and the philosophical possibilities of AI art, without engaging the labor realities that concern practitioners. This pattern mirrors the actor asymmetries Huang and Gadavanij (2025) document in AI-education discourse, where institutional voices systematically overshadow those of affected practitioners.
2. Technical genealogies and spectacle (18.5% of public corpus): These topics focus on novelty, historical lineage, and technological milestones from DeepDream (Mordvintsev et al., 2015) through GANs (Goodfellow et al., 2014) to the Christie’s Belamy auction. This “origin story” discourse captures public attention through spectacle but does not address consent, compensation, or creative integrity.
@@
This pattern is consistent with the idea that the public sphere can operate as an exclusionary space, not through active silencing, but through differential amplification. By flooding discourse with technical, institutional, and spectacular narratives, the environment structurally backgrounds practitioners’ experiential knowledge (Huang and Gadavanij, 2025; Zai et al., 2025; Baeyaert, 2025).
@@ 4.3.2 Algorithmic Agenda-Setting and the “Post-Mediatized” Sphere
A distinctive feature of the contemporary discourse environment is that agenda-setting is increasingly algorithmic. Sichach (2025) argues that attention is now shaped not only by editorial institutions but also by recommender systems and algorithmic ranking that prioritize engagement and scale. Pane (2025) similarly conceptualizes a “post-mediatized” sphere in which algorithmic curation helps set the epistemic boundaries of what is visible and discussable.
@@ 4.4.1 Threat and Utility Travel Well; Governance-Rights Concerns Do Not
One mechanism for differential salience is narrative portability. Media analyses of AI debates show recurring binary frames (e.g., opportunity versus crisis, tool versus threat) that simplify complexity into repeatable tropes (Bøgh, 2025; Cheung, 2024). These become dominant frames—interpretive packages that achieve salience through repetition and institutional amplification, crowding out alternative viewpoints.
@@
By contrast, governance-rights themes (transparency, ownership, compensation) map predominantly to Cluster 17, which receives only 5.84% of public discourse mass despite containing 40.6% of artist probes, a salience ratio of $6.95\times$. Governance-rights themes require technical pipeline details (e.g., how training data is sourced), legal doctrines (e.g., copyright, licensing, consent), and institutional enforcement mechanisms (e.g., opt-out registries, compensation schemes). These elements are more complex to compress into fast-moving media frames and thus more vulnerable to the compression we document (Sichach, 2025; Zai et al., 2025; Pane, 2025).
@@ 4.4.2 Coalition Mobilization Accentuates the Stakes
Recent organizing by creator coalitions underscores that transparency, consent, and compensation are not fringe concerns but actively contested governance claims. Over 100,000 visual artists in the UK have mobilized through the AOI Coalition against unauthorized use of copyrighted works in AI training (AOI Coalition of Artists, 2025); the US Creators Coalition on AI, launched in December 2025, pursues similar objectives (Creators Coalition on AI, 2025). The Andersen v. Stability AI class action (Andersen et al. v. Stability AI et al., 2023) demonstrates that artist grievances have reached federal litigation, with plaintiffs alleging that generative models trained without consent constitute mass infringement.
@@ 4.7 Limitations
Methodological constraints: Artist probes are structured responses constrained by survey framing (Lovato et al., 2024), not organic discourse. The public corpus is shaped by Google Search retrieval, potentially overrepresenting highly visible sources. Our use of a single embedding model (e5-large-v2) may introduce representational biases.
@@ 4.8 Conclusion
Together, these findings constitute an empirical signature of epistemic marginalization (Baeyaert, 2025)—a structural backgrounding of affected-group knowledge that occurs without explicit silencing or malicious intent, through the combined operation of agenda-setting, algorithmic amplification, and milestone-driven spectacle (Sichach, 2025; Pane, 2025; Cheung, 2024).
@@
5 Supplementary Methods
Figure M1: Analysis workflow. Three-stage pipeline: (1) corpus construction and embedding, (2) consensus UMAP projection and clustering, (3) cross-corpus comparison and salience analysis.
5.1 Consensus UMAP Implementation
Our consensus-based projection methodology builds on recent computational approaches for comparing stakeholder discourse in shared semantic spaces (Elmholdt et al., 2025; Matsui and Ferrara, 2024). Below we detail the implementation.
Figure S1: Consensus UMAP methodology. We generate 31 UMAP projections with different random seeds, compute pairwise distance matrices for each, average these matrices, and fit a final embedding from the consensus distance structure. This approach increases average seed-to-consensus Adjusted Rand Index from 0.56 (naive coordinate averaging) to 0.71 (distance-matrix consensus), yielding a more stable reference geometry.
5.1.1 Seed Pool and Multi-Run Generation
  • n_neighbors: 15
  • min_dist: 0.1
  • n_components: 8
  • metric: cosine
5.1.2 Distance-Matrix Consensus
5.1.3 UMAP Hyperparameter Selection
  • min_dist: 0.001 to 0.4 in evenly spaced steps
5.2 Projection Head Architecture
The MLP projection head consists of:
  • Input: 1024-dimensional e5-large-v2 embedding
  • Hidden layers: [512, 256, 128] with ReLU activation and dropout (0.2)
  • Output: 8-dimensional consensus UMAP coordinates
  • Training: Adam optimizer, MSE loss, 80/20 train/validation split
5.3 Robustness Checks
5.3.1 Media Type and Year Associations
  • Media type $\times$ cluster: Cramér’s $V=0.453$
  • Year $\times$ cluster: Cramér’s $V=0.539$
  • Year $\times$ cluster within Articles: Cramér’s $V=0.542$
  • Year $\times$ cluster within Audio: Cramér’s $V=0.433$
  • Year $\times$ cluster within Papers: Cramér’s $V=0.710$
5.4 Use of Large Language Models
LLMs were used for the following tasks:
  1. 1.
    Anchor generation: GPT-5.1 generated 250 synthetic Likert-style statements using a factorial design (5 themes $\times$ 5 agreement levels $\times$ 10 discourse styles; see Table S4)
  2. 2.
    Validation filtering: Human + LLM pipeline (Ziems et al., 2024) retained on-theme public probes
  3. 3.
    Topic interpretation: GPT-5.1 Pro, Gemini Pro 1.5, and Claude Opus 4.5 independently labeled clusters (Ziems et al., 2024); GPT-5.1 Pro served as consolidation judge
Human oversight was maintained throughout, and multiple models were used to mitigate single-model biases (Ziems et al., 2024).
@@ 6 Supplementary Results
Figure S2: Semantic map of all data in two dimensions. Two-dimensional UMAP projection of the public discourse semantic space showing all 22 topic clusters. Topics containing artist perspectives are highlighted. Artist concerns concentrate in a narrow region of the semantic space, while the majority of topics contain no artist perspective.
6.1 Semantic Pathway Analysis
Analysis of per-artist cross-theme patterns reveals:
  • Top 10 semantic pathways account for 89.4% of artists
Figure S3: Semantic pathway concentration analysis. Top 10 pathways account for 89.4% of all artists, indicating stable cross-theme patterns rather than idiosyncratic scattering. Cluster 17 appears in all top 10 pathways for Ownership and Transparency themes; Clusters 10 and 16 capture most Threat and Utility responses.
Figure S4: Theme-to-topic mapping heatmap. Distribution of artist probes (n=1,259n=1{,}259) across public discourse topics by thematic dimension. Affective themes (Threat, Utility) primarily map to Clusters 10 and 16; governance themes (Ownership, Transparency) collapse almost entirely into Cluster 17; compensation themes spread across Clusters 1, 9, and 17. Darker shading indicates higher concentration. All 18 ownership frames—from “Artist Exclusive” to “Anti-Ownership/Commons”—compress into a single topic.
6.2 Base Map Stability Validation
  • Naive coordinate averaging (31 seeds): Mean ARI =0.56=0.56
  • Distance-matrix consensus (31 seeds): Mean ARI =0.71=0.71
  • Relative improvement: $\sim$27%
6.3 Additional Statistical Details
Full permutation test results for geometric separation:
  • Public vs. Artist Probes centroid distance: 2.034 ($p<0.0001$)
  • Public Probes vs. Artist Probes centroid distance: 1.336 ($p<0.0001$)
  • $k$-NN same-source rate (artist probes): 92.1% ($p<0.0001$)
  • $k$-NN same-source rate (public probes vs. artist): 76.5% ($p<0.0001$)
6.4 Search Phrases
Table S1 presents the search phrases used to construct the public discourse corpus.
Table S1: Breadth of search phrases and prevalence of relevant sources. Left: search phrases used to build the public discourse corpus. Right: the number of relevant documents found in the first page of Google search results for each phrase.
@@ 6.4 Search Phrases
6.5 Full Topic Inventory
Table S2 presents the complete 22-topic inventory with labels, keywords, and confidence scores from our multi-annotator interpretation process.
Table S2: Complete list of cluster labels, keywords, and final LLM judge’s confidence in interpretations.
@@ 14 GAN-era AI art explainer/history GANs, Obvious/Belamy (Christie’s), Robbie Barrat, latent space 8
6.6 Cluster Distribution by Data Source
Table S3 shows the distribution of public discourse chunks, public probes, and artist probes across all 22 topics. Bold values indicate notable concentrations.
Table S3: Cluster distribution by data source.
@@ 6.6 Cluster Distribution by Data Source
6.7 LLM-Generated Likert Anchor Statements
Table S4 presents representative examples from the 250 synthetic Likert-style anchor statements generated by GPT-5.1, organized by theme and agreement level. Each cell in the 5 $\times$ 5 design (theme $\times$ Likert level) was populated with 10 discourse-style variants drawn from genres present in our public corpus (blog opinion, news editorial, artist interview, legal brief, policy report, panel discussion, first-person narrative, union statement, corporate statement, academic article, etc.). The full set of 250 anchor statements is available as likert_thematic_phrases.csv in our code repository.
Table S4: Design matrix and representative examples of the 250 LLM-generated Likert-style anchor statements (5 themes $\times$ 5 agreement levels $\times$ 10 discourse styles = 250). One example per theme–level combination is shown; the full set is available in our code repository.
Theme Level Representative Example (discourse style)
Utility Str. disagree “From my perspective, artificial intelligence art models are ruining the field of art rather than improving it.” (blog opinion)
Utility Disagree “The panel expressed skepticism that artificial intelligence art models truly benefit the field of art.” (panel discussion)
Utility Neutral “The report takes no position on whether artificial intelligence art models are good or bad for art.” (research report)
Utility Agree “The article argues that artificial intelligence art models open exciting new possibilities for artists.” (news opinion)
Utility Str. agree “This manifesto celebrates artificial intelligence art models as a transformative good for creative practice.” (manifesto)
Ownership Str. disagree “In my view, artificial intelligence generated artwork in an artist’s style should not be considered the artist’s property at all.” (legal opinion)
Ownership Disagree “This blog post argues that users should have primary control over artificial intelligence generated images, rather than the original artists.” (blog opinion)
Ownership Neutral “The article maps the legal uncertainties around ownership of artificial intelligence generated images in an artist’s style.” (academic article)
Ownership Agree “The union calls for recognizing artificial intelligence generated artworks in an artist’s style as that artist’s property.” (union statement)
Ownership Str. agree “The manifesto declares that artificial intelligence generated images in an artist’s style belong entirely to that artist.” (manifesto)
Transparency Str. disagree “Our company rejects proposals to mandate detailed transparency about the art and images used to train our artificial intelligence systems.” (corporate statement)
Transparency Disagree “The brief raises concerns about broad requirements to reveal all art and images used to train artificial intelligence systems.” (policy brief)
Transparency Neutral “Survey respondents were split on whether artificial intelligence model creators should be required to disclose training data details.” (survey report)
Transparency Agree “In the interview, the illustrator says that companies should have to tell us which collections they used to train their artificial intelligence models.” (artist interview)
Transparency Str. agree “The report strongly recommends strict disclosure mandates covering all art and images used to train artificial intelligence models.” (policy report)
Threat Str. disagree “The paper contends that artificial intelligence art models complement, rather than threaten, human artistic labor.” (academic article)
Threat Disagree “In this interview, the painter says they are not overly worried about artificial intelligence art systems taking their job.” (artist interview)
Threat Neutral “The news story quotes some artists who fear displacement by artificial intelligence and others who do not.” (news report)
Threat Agree “Survey data show that many art workers view artificial intelligence art tools as a significant career risk.” (survey report)
Threat Str. agree “Our statement declares that artificial intelligence art models are a direct attack on working artists and their jobs.” (union statement)
Compensation Str. disagree “The report concludes that mandatory artist compensation for artificial intelligence training would be inappropriate.” (policy report)
Compensation Disagree “I am not convinced that artists deserve extensive compensation every time their work informs artificial intelligence art models.” (first person)
Compensation Neutral “Our panel debated revenue sharing, lump sum payments, and public funds for artists affected by artificial intelligence art.” (panel discussion)
Compensation Agree “The paper proposes a licensing model in which artists are paid when their art appears in artificial intelligence training sets.” (academic article)
Compensation Str. agree “This manifesto demands strong profit sharing and royalty rights for artists in all artificial intelligence art contexts.” (manifesto)